OpenStudy (anonymous):

If x*x + 2x - 35 = 0, then x = ?

OpenStudy (anonymous):

is that x^2

OpenStudy (tyteen4a03):

@Kanwar245 Yes. Remember our quadratic formula: \(x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\)

OpenStudy (anonymous):

what is x??

OpenStudy (tyteen4a03):

@HAPPY10 Substitute the coefficients of the quadratic equation into the quadratic formula, and you will get the solution(s) to x.

OpenStudy (radar):

Factor it unless you are told to use quadratic formulae. (x+7)(x-5)=0

OpenStudy (anonymous):

thanks